OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step?

OpenStudy (anonymous):

2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 2x + (4 - x) = 8 2x + y = x + y 3x + 2y = 12 x = 4

OpenStudy (anonymous):

I think it is 2x+ (4-x) = 8

OpenStudy (anonymous):

i just need it checked

OpenStudy (anonymous):

I'll give a medal!

OpenStudy (aripotta):

wait, what are the choices? i'm confused

OpenStudy (anonymous):

The choices are 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 2x + (4 - x) = 8 2x + y = x + y 3x + 2y = 12 x = 4 I know it isn't the last one... I think It is 2x + (4 - x) = 8 I just need it checked

OpenStudy (aripotta):

it's in order, isn't it?

OpenStudy (aripotta):

oh idk. sorreh.

OpenStudy (anonymous):

I think so...but IDK! I SO CONFUSED! I hate these problems

OpenStudy (anonymous):

Its ok... Idont know either! :( HATE MATH RIGHT NOW!!

OpenStudy (aripotta):

the first step should be (x + y = 4)(-1)---> -x - y = -4 to eliminate the y...

OpenStudy (anonymous):

Yea, it makes since... but I dont know it they mean rewritng the question as the first step... IDK

OpenStudy (aripotta):

i'd go with the second choice

OpenStudy (aripotta):

but that's just me

OpenStudy (anonymous):

Ok thanks so much! AriPotta... I will give you an award! :) thanks so much!

OpenStudy (aripotta):

no problem :)